OpenStudy (anonymous):

Does |x + 6| - 5 = -2 have no solution

hero (hero):

Hmm, let's see: Add 5 to both sides to get: |x + 6| = -2 + 5 |x + 6| = 3 x + 6 = 3 or x + 6 = -3 looks like it has two solutions. All you have to do now is finish solving for x.

OpenStudy (anonymous):

alright thanks